### v1

@@ 6.4. Dimension increment and bootstrap
We combine the results of the three previous subsections to show that the dimensional properties of a prescribed measure $\nu$ on $X$ are improved under the action of the $\mu$-random walk on $X$. This is Proposition˜6.16. By iteration, we deduce the desired bootstrap to high dimension, Proposition˜6.2.
@@
Let $\kappa,\varepsilon,\rho\in(0,1/10)$, $\alpha\in{[\kappa,1-\kappa]}$, $\tau\geq 0$ be some parameters. Consider on $X$ a Borel measure $\nu$ of mass at most $1$ and which is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},\tau)$-robust. Denote by $n_{\rho}\geq 0$ the integer part of $\frac{1}{10\ell}\lvert\log\rho\rvert$.
@@
We may assume $\tau=0$. We write $n=n_{\rho}$. By Proposition˜4.1 integrated over $\nu$, we have
$\mu^{*n}*\nu\{\operatorname{inj}\leq\rho^{1/2}\}\ll\rho^{c/2}(e^{-cn}\rho^{-C\varepsilon}+1)$
@@
(6.23) $\sup_{y\in X}\nu^{\prime}(B_{\rho^{1/2}}y)\leq\rho^{\frac{1}{2}(\alpha+\varepsilon)\dim X}.$
To this end, we first linearize the situation by looking through the covering of charts from §6.3. More precisely, we apply Lemma˜6.15 with parameters $\eta=\rho^{\varepsilon}$ and $\delta=\rho^{1/3}$. This yields a map $\varphi:\{\operatorname{inj}\geq\rho^{\varepsilon}\}\rightarrow B^{\mathfrak{g}}_{1}$, we set ${\tilde{\nu}}=\varphi_{\star}\nu$. The assumption that $\nu$ is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},0)$-robust and has mass at most $1$ implies, via Lemma˜6.15 item 1) and provided $\rho\lll_{\varepsilon}1$, that for every $r\geq\rho$,
@@
We now aim to apply Proposition˜6.14 to the measure $\tilde{\nu}$}, and for the random box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$ where $g\sim\mu^{*n}$, or rather its close companion
@@
which is a good approximation of $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$ (by (6.25) below), and whose partial flag we know how to control thanks to §6.1. Indeed, by Proposition˜6.4 and Corollary˜6.5, the distributions of $(\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-})_{g\sim\mu^{*n}}$ and $(\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0})_{g\sim\mu^{*n}}$ satisfy respectively $\operatorname{(MNC)}$ and $\operatorname{(MNC)}^{\perp}$ with parameters $(e^{-n},C,c)$, or equivalently $(\rho,C,c)$ up to dividing $c$ by $11\ell$. Here $C,c>0$ are constants that only depend on $\mu$.
Provided $\rho,\varepsilon\lll 1$, the multislicing Proposition˜6.14 yields a subset $E_{1}\subseteq G$ and some constant $\varepsilon_{0}=\varepsilon_{0}(\mu)>0$ such that $\mu^{*n}(E_{1})\leq\rho^{\varepsilon_{0}}$ and for $g\in G\smallsetminus E_{1}$, there exists a set $\tilde{A}_{g}\subseteq\mathfrak{g}$ with ${\tilde{\nu}}(\mathfrak{g}\smallsetminus\tilde{A}_{g})\leq\rho^{\varepsilon_{0}}$ and such that for every $v\in\mathfrak{g}$,
(6.24) ${\tilde{\nu}}_{|\tilde{A}_{g}}\left(B^{\mathscr{V}_{g}}_{\rho^{\bf t}}+v\right)\leq\rho^{\varepsilon_{0}}\operatorname{Leb}\left(B^{\mathscr{V}}_{\rho^{\bf t}}\right)^{\alpha}.$
On the other hand, by the large deviation principle for $\log\mathtt{r}_{g}$ and Lemma˜3.5 item (i), there exists a subset $E_{2}\subseteq G$ and a constant $\gamma=\gamma(\mu,\varepsilon)>0$ such that $\mu^{*n}(E_{2})\ll\rho^{\gamma}$, and for every $g\in G\smallsetminus E_{2}$, we have $\mathtt{r}_{g}\in[\rho^{\frac{1}{10}+\varepsilon},\rho^{\frac{1}{10}-\varepsilon}]$ and $\|\mathtt{b}_{g}\|\leq\rho^{-\varepsilon}$. Combined with (6.3), we obtain
(6.25) $B^{\mathscr{V}_{g}}_{\rho^{\bf t}}\,\overset{\rho^{-O(\varepsilon)}}{\simeq}\,\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}.$
@@
(6.26) ${\tilde{\nu}}_{|\tilde{A}_{g}}\left(\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}+v\right)\leq\rho^{\varepsilon_{0}-O(\varepsilon)}\operatorname{Leb}\left(B^{\mathscr{V}}_{\rho^{\bf t}}\right)^{\alpha}.$
We now get back to $X$. To control the distortion of $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$, we observe for $g\in G\smallsetminus E_{2}$, we have $B^{\mathfrak{g}}_{\rho^{2/3}}\subseteq\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}\subseteq B^{\mathfrak{g}}_{\rho^{1/3}}$ provided $\varepsilon\lll 1$. Applying Lemma˜6.15 item $2)$ and (6.26), we deduce that for all $g\in G\smallsetminus\cup_{i=1}^{2}E_{i}$, setting $A_{g}=\varphi^{-1}(\tilde{A}_{g})$, we have $\nu(G\smallsetminus A_{g})\leq\rho^{\varepsilon_{0}}$ and for every $y\in G$,
@@
Taking $\nu^{\prime}=\int_{G\smallsetminus\cup_{i=1}^{2}E_{i}}g_{\star}\nu_{|A_{g}}\,\mathrm{d}\mu^{*n}(g)$, and $\varepsilon\lll\varepsilon_{0}$, $\rho\lll_{\varepsilon}1$, this concludes the proof of (6.23), whence that of the proposition. ∎
We now deduce high dimension (Proposition˜6.2) from the combination of effective recurrence (Proposition˜4.1), initial positive dimension (Proposition˜5.1), and dimension increment (Proposition˜6.16).
Proof of Proposition˜6.2.
Let $A>0$ be a large enough constant depending on the initial data $\mu$. Combining Proposition˜5.1 and Proposition˜4.1, we may assume $\kappa>0$ small enough from the start, so that for any $M>0$, for every $\rho\lll_{M}1$ and $n\geq M\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}(x)\rvert$, the measure
@@
By Proposition˜6.16, there is some small constant $\varepsilon=\varepsilon(\mu,\kappa)>0$, such that up to imposing from the start $M\ggg_{\kappa}1$, we have for all $\rho\lll_{\kappa,M}1$, all $n\geq(\frac{1}{10\ell}+1)M\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}(x)\rvert$ and $r\in[\rho^{2M\varepsilon},\rho^{1/(2M\varepsilon)}]$,
@@
These estimates for single scales can be combined using [6, Lemma 4.5] to get under the same conditions:
$\text{{$\mu^{*n}*\delta_{x}$} is $(\kappa+\varepsilon,{\mathcal{B}}_{[\rho^{M\varepsilon},\rho^{1/(M\varepsilon)}]},O_{\kappa,M}(\rho^{\kappa/M}))$-robust}.$
The argument in the last paragraph can be applied iteratively, adding at each step $k$ the value $\varepsilon$ to the dimension provided the latter is not yet above $1-\kappa$ and provided $M\ggg_{\kappa,k}1$. As the value of $\varepsilon$ only depends on $\mu,\kappa$, we reach dimension $1-\kappa$ in a finite number of steps. This concludes the proof. ∎
@@ 7. From high dimension to equidistribution
In this section, we establish ˜1.2 and ˜1.3. We further establish a double equidistribution estimate (Proposition˜7.3) which will be useful to prove the divergent case of ˜1.1.
@@
(7.1) $|\eta_{t}*\nu(f)|\leq(\rho^{\kappa}+\tau){\mathcal{S}}_{\infty,l}(f).$
@@
The proof is similar to that of [7, Proposition 5.1]. We provide a sketch for completeness and refer the reader to [7] for details.
@@
(7.2) $\|P_{\eta_{t}}f\|_{L^{2}}\ll t^{-c}{\mathcal{S}}_{2,l}(f).$
The proof of (7.2) exploits the quantitative decay of matrix coefficients (see [2, Lemma 3] and [22, Equations (6.1), (6.9)]):
$\exists\delta_{0}=\delta_{0}(\Lambda)>0,\,\forall g\in G,\quad\lvert\langle f(g\,\cdot),f\rangle_{L^{2}}\rvert\ll\lVert g\rVert^{-\delta_{0}}{\mathcal{S}}_{2,l}(f)^{2},$
and the non-concentration property of $\sigma$ from Proposition˜3.2, see [7, Proposition 5.2] for details.
@@
The first integral in the right hand side is bounded by $\rho\operatorname{Lip}(P_{\eta_{t}}f)\ll\rho t{\mathcal{S}}_{\infty,l}(f).$ To bound the second integral, note we may assume from the start $\tau=0$. Then $\nu_{\rho}$ satisfies $\mathrm{d}\nu_{\rho}(x)=\frac{\nu(B_{\rho}x)}{m_{G}(B_{\rho})}\,\mathrm{d}m_{X}(x)\ll\rho^{-\kappa\dim X}\,\mathrm{d}m_{X}(x)$. Applying (7.2), we find the second term in the right hand side is bounded by $O(t^{-c}\rho^{-\kappa\dim X}{\mathcal{S}}_{2,l}(f)).$ The proof is concluded by taking $t\in[\rho^{-1/2},\rho^{-1/4}]$, $\kappa$ small enough in terms of $c,\dim X$, and $\rho_{0}$ small enough in terms of $G$, $\kappa$. ∎
@@
We are now able to conclude the proof of ˜1.2. The strategy is to use Lemma˜7.2 to decompose $\eta_{t}$ as a random walk part $\mu^{*n}$ (where $n=n(\mu,t)$) which generates high dimension thanks to Proposition˜6.2, followed by some $\eta_{t^{\prime}}$-part (with $t^{\prime}=t^{\prime}(\mu,t)$) which will convert this high dimension into equidistribution via Proposition˜7.1. The apparent obstruction is that the decomposition appearing Lemma˜7.2 does not separate the $\mu$ part and the $\eta$ part, because the term $\delta_{k_{g}}*\eta_{t\mathtt{r}_{g}}$ involves $g$. To deal with this obstacle, we partition the space of parameters $g$ into $O(\rho^{-\alpha})$ subsets ($\alpha\lll_{\kappa}1$) in which $\delta_{k_{g}}*\eta_{t\mathtt{r}_{g}}$ hardly depends on $g$.
@@
Up to replacing $\lambda$ by a suitable convolution power until a stopping time (as in [7, Lemma 5.3]), we may assume that $\lambda$-almost every $\phi$ is orientation preserving. This way we place ourselves in the context of Section˜2, and all results obtained until now are applicable.
@@
(7.3) $\lvert\delta_{k_{0}}*\eta_{tr_{0}}*\nu(f)-\delta_{k_{1}}*\eta_{tr_{1}}*\nu(f)\rvert\ll(\lVert\operatorname{Id}-k_{0}k_{1}^{-1}\rVert+\lvert\log r_{0}r_{1}^{-1}\rvert){\mathcal{S}}_{\infty,l}(f)\nu(X).$
@@
$P^{\prime}=\bigsqcup_{\mathtt{r}\in{\mathscr{R}},i\in I}P^{\prime}(\mathtt{r},i).$
Hence, by Lemma˜7.2 and (7.3), we obtain for every $n\geq 1$,
@@
(7.4) $\leq\sum_{i\in I,r\in{\mathscr{R}}}\lvert\delta_{k_{i}}*\eta_{tr}*\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}(f)\rvert+O(\rho^{\alpha}{\mathcal{S}}_{\infty,l}(f)).$
We now bound each term in the sum from (7). Let $\kappa=\kappa(\Lambda,\mu)>0$ as in Proposition˜7.1. Assume $\operatorname{inj}(x)\geq\rho$. By Proposition˜6.2, there are constants $C=C(\Lambda,\mu)>1$ and $\varepsilon_{1}=\varepsilon_{1}(\Lambda,\mu)>0$ such that, provided $\rho\lll 1$, the measure $\mu^{*n}*\delta_{x}$ on $X$ is $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust for any $n\geq C\lvert\log\rho\rvert$. For the rest of this proof, we choose $\rho$ and $n$ depending on $\Lambda,\mu,t$ so that
(7.5) $t=\rho^{-C\ell-3/8}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,n=\lceil C\lvert\log\rho\rvert\rceil$,
@@
(7.6) $\mu^{*n}\{g\,:\,\mathtt{r}_{g}\notin{\mathscr{R}}^{\prime}\}\leq\rho^{\varepsilon_{2}}$
On the other hand, for $i\in I$, $\mathtt{r}\in{\mathscr{R}}^{\prime}$, observing that $\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}\leq\mu^{*n}*\delta_{x}$ is $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust, and ${\mathcal{S}}_{\infty,l}(f\circ k_{i})\ll{\mathcal{S}}_{\infty,l}(f)$ , $m_{X}(f\circ k_{i})=m_{X}(f)$, we get via Proposition˜7.1, for $t\ggg 1$,
(7.7) $\lvert\eta_{tr}*\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}(f)\rvert\leq(\rho^{\kappa}+\rho^{\varepsilon_{1}}){\mathcal{S}}_{\infty,l}(f).$
Combining (7.5), (7.6), (7.7) and choosing $\alpha$ small enough in terms of $\varepsilon_{1},\varepsilon_{2},\kappa$, we obtain the bound announced by ˜1.2. So far, we have worked under the condition $\operatorname{inj}(x)\geq t^{-(C\ell+3/8)^{-1}}$. Noting the claim is trivial otherwise, the proof of ˜1.2 is complete.
@@
By a similar argument, we see Proposition˜7.1 is still valid with $(t,\eta_{t})$ replaced by $(e^{k},\mu^{k})$ Replacing (7.2) by the equality $\mu^{*(k+n)}=\mu^{*k}*\mu^{*n}$, we can then argue as in the proof of ˜1.2. Details are left to the reader. ∎
@@
(7.8) $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2}):=\left\lvert\int_{\mathbb{R}^{d}}f_{1}\bigl{(}a(t_{1})u(\boldsymbol{s})x_{0}\bigr{)}f_{2}\bigl{(}a(t_{2})u(\boldsymbol{s})x_{0}\bigr{)}\,\mathrm{d}\sigma(\boldsymbol{s})-m_{X}(f_{1})m_{X}(f_{2})\right\rvert$.
@@
(7.9) $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}.$
@@
The proof is the same as that of [7, Proposition 6.1], using ˜1.2 in the place of [7, Theorem B’]. ∎

### v2

@@ 7.4. Dimension increment and bootstrap
We combine the results of the three previous subsections to show that the dimensional properties of a prescribed measure $\nu$ on $X$ are improved under the action of the $\mu$-random walk on $X$. This is Proposition˜7.16. By iteration, we deduce the desired bootstrap to high dimension, Proposition˜7.2.
@@
Let $\kappa,\varepsilon,\rho\in(0,1/10)$, $\alpha\in{[\kappa,1-\kappa]}$, $\tau\geq 0$ be some parameters. Consider on $X$ a Borel measure $\nu$ of mass at most $1$, which is $(\alpha,{\mathcal{B}}_{[\rho^{2/3},\rho^{1/3}]},\tau)$-robust, and supported on $\{\operatorname{inj}\geq\delta^{\varepsilon}\}$. Denote by $n_{\rho}\geq 0$ the integer part of $\frac{1}{10\ell}\lvert\log\rho\rvert$.
@@
We may assume $\tau=0$, and $\rho$ small enough in terms of $\varepsilon$ as well (not only $\Lambda,\mu,\kappa$). We write $n=n_{\rho}$. By Proposition˜5.1 integrated over $\nu$, we have
$\mu^{*n}*\nu\bigl\{\operatorname{inj}\leq\rho^{1/2}\bigr\}\ll\rho^{c/2}(e^{-cn}\rho^{-C\varepsilon}+1)$
@@
$\sup_{y\in X}\nu^{\prime}\bigl(B_{\rho^{1/2}}y\bigr)\leq\rho^{\frac{1}{2}(\alpha+\varepsilon)\dim X}.$ (7.31)
To this end, we first linearize the situation by looking through the covering of charts from §7.3. More precisely, we apply Lemma˜7.15 with parameter $\delta=\rho^{1/3}$. This yields a map $\varphi:\{\operatorname{inj}\geq\rho^{1/3}\}\rightarrow B^{\mathfrak{g}}_{1}$, we set ${\tilde{\nu}}=\varphi_{\star}\nu$. The assumption that $\nu$ is $(\alpha,{\mathcal{B}}_{[\rho^{2/3},\rho^{1/3}]},0)$-robust and has mass at most $1$ implies, via Lemma˜7.15 item 1) and provided $\rho\lll_{\varepsilon}1$, that for every $r\in[\rho^{2/3},\rho^{1/3}]$,
@@
We now aim to apply Proposition˜7.14 to the measure $\tilde{\nu}$}, and for the random box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$ where $g\sim\mu^{*n}$, or rather its close companion
@@
which is a good approximation of $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$ (by (7.33) below), and whose partial flag we know how to control thanks to §7.1. Indeed, by Proposition˜7.4 and Corollary˜7.5, the distributions of $(\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-})_{g\sim\mu^{*n}}$ and $(\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0})_{g\sim\mu^{*n}}$ satisfy respectively $\operatorname{(MNC)}$ and $\operatorname{(MNC)}^{\perp}$ with parameters $(e^{-n},C,c)$, or equivalently $(\rho,C,c)$ up to dividing $c$ by $11\ell$. Here $C,c>0$ are constants that only depend on $\mu$.
Provided $\rho,\varepsilon\lll 1$, the multislicing Proposition˜7.14 yields a subset $E_{1}\subseteq G$ and some constant $\varepsilon_{0}=\varepsilon_{0}(\mu)>0$ such that $\mu^{*n}(E_{1})\leq\rho^{\varepsilon_{0}}$ and for $g\in G\smallsetminus E_{1}$, there exists a set $\tilde{A}_{g}\subseteq\mathfrak{g}$ with ${\tilde{\nu}}(\mathfrak{g}\smallsetminus\tilde{A}_{g})\leq\rho^{\varepsilon_{0}}$ and such that for every $v\in\mathfrak{g}$,
${\tilde{\nu}}_{|\tilde{A}_{g}}\left(B^{\mathscr{V}_{g}}_{\rho^{\bf t}}+v\right)\leq\rho^{\varepsilon_{0}}\operatorname{Leb}\left(B^{\mathscr{V}}_{\rho^{\bf t}}\right)^{\alpha}.$ (7.32)
On the other hand, by the large deviation principle for $\log\mathtt{r}_{g}$ and Lemma˜3.3, there exists a subset $E_{2}\subseteq G$ and a constant $\gamma=\gamma(\mu,\varepsilon)>0$ such that $\mu^{*n}(E_{2})\ll\rho^{\gamma}$, and for every $g\in G\smallsetminus E_{2}$, we have $\mathtt{r}_{g}\in[\rho^{\frac{1}{10}+\varepsilon},\rho^{\frac{1}{10}-\varepsilon}]$ and $\|\mathtt{b}_{g}\|\leq\rho^{-\varepsilon}$. In view of (7.3), we obtain in particular
$B^{\mathscr{V}_{g}}_{\rho^{\bf t}}\,\overset{\rho^{-O(\varepsilon)}}{\simeq}\,\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}.$ (7.33)
@@
${\tilde{\nu}}_{|\tilde{A}_{g}}\left(\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}+v\right)\leq\rho^{\varepsilon_{0}-O(\varepsilon)}\operatorname{Leb}\left(B^{\mathscr{V}}_{\rho^{\bf t}}\right)^{\alpha}.$ (7.34)
We now get back to $X$. To control the distortion of $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$, we observe for $g\in G\smallsetminus E_{2}$, we have $B^{\mathfrak{g}}_{\rho^{2/3}}\subseteq\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}\subseteq B^{\mathfrak{g}}_{\rho^{1/3}}$ provided $\varepsilon\lll 1$. Applying Lemma˜7.15 item $2)$ and (7.34), we deduce that for all $g\in G\smallsetminus(E_{1}\cup E_{2})$, setting $A_{g}=\varphi^{-1}(\tilde{A}_{g})$, we have $\nu(G\smallsetminus A_{g})\leq\rho^{\varepsilon_{0}}$ and for every $y\in X$,
@@
Taking $\nu^{\prime}=\int_{G\smallsetminus(E_{1}\cup E_{2})}g_{\star}\nu_{|A_{g}}\,\mathrm{d}\mu^{*n}(g)$, and $\varepsilon\lll\varepsilon_{0}$, $\rho\lll_{\varepsilon}1$, this concludes the proof of (7.31), whence that of the proposition. ∎
We now deduce high dimension (Proposition˜7.2) from the combination of effective recurrence (Proposition˜5.1), initial positive dimension (Proposition˜6.1), and dimension increment (Proposition˜7.16).
Proof of Proposition˜7.2.
Let $A>0$ be a large enough constant depending on the initial data $\mu$. Combining Proposition˜6.1 and Proposition˜5.1, we may assume $\kappa>0$ small enough from the start, so that for any $M>0$, for every $\rho\lll_{M}1$ and $n\geq M\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}(x)\rvert$, the measure
@@
By Proposition˜7.16, there is some small constant $\varepsilon=\varepsilon(\mu,\kappa)>0$, such that up to imposing from the start $M\ggg_{\kappa}1$, we have for all $\rho\lll_{\kappa,M}1$, all $n\geq(\frac{1}{10\ell}+1)M\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}(x)\rvert$ and $r\in[\rho^{3M/2},\rho^{3/M}]$,
@@
These estimates for single scales can be combined using [4, Lemma 4.5] to get under the same conditions:
$\text{{$\mu^{*n}*\delta_{x}$} is $(\kappa+\varepsilon,{\mathcal{B}}_{[\rho^{3M/4},\rho^{3/(2M)}]},O_{\kappa,M}(\rho^{\kappa/M}))$-robust}.$
The argument in the last paragraph can be applied iteratively, adding at each step the value $\varepsilon$ to the dimension provided that the latter is not yet above $1-\kappa$. Noting the value of $\varepsilon$ only depends on $\mu,\kappa$, we reach dimension $1-\kappa$ in at most $m:=\lceil\varepsilon^{-1}\rceil$ steps. Iteration is allowed provided that $M$ is chosen large enough to satisfy $(\frac{3}{4})^{m}M>(\frac{3}{2})^{m}\frac{1}{M}$, i.e. $M>2^{m/2}$. We could further require $M>(3/2)^{m}$ from the beginning to guarantee that $\rho$ belongs to the range of scales for which we have $(1-\kappa)$-robustness when the iteration ends. This concludes the proof. ∎
@@ 8. From high dimension to equidistribution
In this section, we establish ˜1.2 and ˜1.3. We further establish a double equidistribution estimate (Proposition˜8.3) which will be useful to prove the divergent case of ˜1.1.
@@
$|\eta_{t}*\nu(f)|\leq(\rho^{\kappa}+\tau){\mathcal{S}}_{\infty,l}(f).$ (8.1)
@@
The proof is similar to that of [3, Proposition 5.1]. We provide a sketch for completeness and refer the reader to [3] for details.
@@
$\|P_{\eta_{t}}f\|_{L^{2}}\ll t^{-c}{\mathcal{S}}_{2,l}(f).$ (8.2)
The proof of (8.2) exploits the quantitative decay of matrix coefficients (see [2, Lemma 3] and [23, Equations (6.1), (6.9)]):
$\exists\delta_{0}=\delta_{0}(\Lambda)>0,\,\forall g\in G,\qquad\lvert\langle f(g\,\cdot),f\rangle_{L^{2}}\rvert\ll\lVert g\rVert^{-\delta_{0}}{\mathcal{S}}_{2,l}(f)^{2},$
and the non-concentration property of $\sigma$ from Proposition˜4.2, see [3, Proposition 5.2] for details.
@@
The first integral in the right hand side is bounded by $\rho{\mathcal{S}}_{\infty,1}(P_{\eta_{t}}f)\ll\rho t{\mathcal{S}}_{\infty,1}(f)$. To bound the second integral, note we may assume from the start $\tau=0$. Then $\nu_{\rho}$ satisfies $\mathrm{d}\nu_{\rho}(x)=\frac{\nu(B_{\rho}x)}{m_{G}(B_{\rho})}\,\mathrm{d}m_{X}(x)\ll\rho^{-\kappa\dim X}\,\mathrm{d}m_{X}(x)$. Applying (8.2), we find the second term in the right hand side is bounded by $O(t^{-c}\rho^{-\kappa\dim X}{\mathcal{S}}_{2,l}(f)).$ The proof is concluded by taking $t\in[\rho^{-1/2},\rho^{-1/4}]$, $\kappa$ small enough in terms of $c,\dim X$, and $\rho_{0}$ small enough in terms of $G$, $\kappa$. ∎
@@
We are now able to conclude the proof of ˜1.2. The strategy is to use Lemma˜8.2 to decompose $\eta_{t}$ as a random walk part $\mu^{*n}$ (where $n=n(\mu,t)$) which generates high dimension thanks to Proposition˜7.2, followed by some $\eta_{t^{\prime}}$-part (with $t^{\prime}=t^{\prime}(\mu,t)$) which will convert this high dimension into equidistribution via Proposition˜8.1. The apparent obstruction is that the decomposition appearing Lemma˜8.2 does not separate the $\mu$ part and the $\eta$ part, because the term $\delta_{k_{g}}*\eta_{t\mathtt{r}_{g}}$ involves $g$. To deal with this obstacle, we partition the space of parameters $g$ into $O(\rho^{-\alpha})$ subsets ($\alpha\lll_{\kappa}1$) in which $\delta_{k_{g}}*\eta_{t\mathtt{r}_{g}}$ hardly depends on $g$.
@@
Up to replacing $\lambda$ by a suitable convolution power until a stopping time (as in [3, Lemma 5.3]), we may assume that $\lambda$-almost every $\phi$ is orientation preserving. This way we place ourselves in the context of Section˜2, and all results obtained until now are applicable.
@@
$\lvert\delta_{k_{0}}*\eta_{tr_{0}}*\nu(f)-\delta_{k_{1}}*\eta_{tr_{1}}*\nu(f)\rvert\ll(\lVert\operatorname{Id}-k_{0}k_{1}^{-1}\rVert+\lvert\log r_{0}r_{1}^{-1}\rvert){\mathcal{S}}_{\infty,l}(f)\nu(X).$ (8.3)
@@
$P^{\prime}=\bigsqcup_{r\in{\mathscr{R}},i\in I}P^{\prime}(i,r).$
Hence, by Lemma˜8.2 and (8.3), we obtain for every $n\geq 1$,
@@
$\leq\sum_{i\in I,r\in{\mathscr{R}}}\lvert\delta_{k_{i}}*\eta_{tr}*\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}(f)\rvert+O(\rho^{\alpha}{\mathcal{S}}_{\infty,l}(f)).$ (8.4)
We now bound each term in the sum from (8). Let $\kappa=\kappa(\Lambda,\mu)>0$ as in Proposition˜8.1. Assume $\operatorname{inj}(x)\geq\rho$. By Proposition˜7.2, there are constants $C=C(\Lambda,\mu)>1$ and $\varepsilon_{1}=\varepsilon_{1}(\Lambda,\mu)>0$ such that, provided $\rho\lll 1$, the measure $\mu^{*n}*\delta_{x}$ on $X$ is $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust for any $n\geq C\lvert\log\rho\rvert$. For the rest of this proof, we choose $\rho$ and $n$ depending on $\Lambda,\mu,t$ so that
$t=\rho^{-C\ell-3/8}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,n=\lceil C\lvert\log\rho\rvert\rceil$, (8.5)
@@
$\mu^{*n}\{g\,:\,\mathtt{r}_{g}\notin{\mathscr{R}}^{\prime}\}\leq\rho^{\varepsilon_{2}}$ (8.6)
On the other hand, for $i\in I$, $r\in{\mathscr{R}}^{\prime}$, observing that $\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}\leq\mu^{*n}*\delta_{x}$ is $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust, and ${\mathcal{S}}_{\infty,l}(f\circ k_{i})\ll{\mathcal{S}}_{\infty,l}(f)$ , $m_{X}(f\circ k_{i})=m_{X}(f)$, we get via Proposition˜8.1, for $t\ggg 1$,
$\lvert\eta_{tr}*\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}(f)\rvert\leq(\rho^{\kappa}+\rho^{\varepsilon_{1}}){\mathcal{S}}_{\infty,l}(f).$ (8.7)
Combining (8.5), (8.6), (8.7) and choosing $\alpha$ small enough in terms of $\varepsilon_{1},\varepsilon_{2},\kappa$, we obtain the bound announced by ˜1.2. So far, we have worked under the condition $\operatorname{inj}(x)\geq t^{-(C\ell+3/8)^{-1}}$. Noting the claim in ˜1.2 is trivial otherwise, the proof is complete.
@@
By a similar argument, we see Proposition˜8.1 is still valid with $(t,\eta_{t})$ replaced by $(e^{k},\mu^{k})$. Replacing (8.2) by the equality $\mu^{*(k+n)}=\mu^{*k}*\mu^{*n}$, we can then argue as in the proof of ˜1.2. Details are left to the reader. ∎
@@
$\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2}):=\left\lvert\int_{\mathbb{R}^{d}}f_{1}\bigl(a(t_{1})u(\boldsymbol{s})x_{0}\bigr)f_{2}\bigl(a(t_{2})u(\boldsymbol{s})x_{0}\bigr)\,\mathrm{d}\sigma(\boldsymbol{s})-m_{X}(f_{1})m_{X}(f_{2})\right\rvert$. (8.8)
@@
$\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}.$ (8.9)
@@
The proof is the same as that of [3, Proposition 6.1], using ˜1.2 in the place of [3, Theorem B’]. ∎